A random sample of 40 households has an average water consumption of 29 cubic meters per month with a standard deviation of 8 cubic meters. Give the 90% confidence interval for the mean usage of water per month.


1
Expert's answer
2022-04-07T13:31:14-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{\sigma}{\sqrt{n}},"

where:

  • "\bar{x}=29" : sample mean
  • z:  the chosen z-value, for a 90% confidence interval z = 1.645
  • "\sigma=8" :  sample standard deviation
  • n = 40 sample size.

So,

"CI=29 \pm 1.645\cdot\cfrac{8}{\sqrt{40}}=\\
=29\pm2. 08=(26.92, 31.08)."